OpenStudy (anonymous):

There is a sales tax $14 of on an item that costs $147 before tax. A second item costs 178.50 before tax. What is the sales tax on the second item?

OpenStudy (thurscon):

I'm assuming that the rate of the sales tax is the same, since that information is not explicitly given in the word problem. We would solve this problem with proportions - \[14/147 = x/178.50\] Then, we would multiply cross products ("property of proportions") \[14(178.50) = 147(x)\] This simplifies to \[2499=147(x)\] Now, we want to isolate the x so that we can find our answer. We do this by dividing both sides by 147. \[2499/147 = 147(x)/147\] So we are left with \[17 = x\] which is your final answer.

OpenStudy (anonymous):

thank you

OpenStudy (thurscon):

Sure, any time! :)